OpenStudy (babybritt002):

The population of a certain single-cell organism doubles every hour, starting with 100 cells. Which COULD be an equation to model this situation, where f(t) is the number of organisms and t is time in hours?

OpenStudy (ranga):

Exponential equations have the general form: f(t) = A * (B)^t when t = 0, f(t) = 100 100 = A * B^0 = A f(t) = 100*(B)^t when t = 1, the population doubles to 200 200 = 100 * B^1 B = 200/100 = 2 Therefore, f(t) = 100 * (2)^t